I'm trying to figure out how to find the square root of 49x^8 and simplify
Answer by stanbon(75887)

I'm trying to figure out how to find the square root of 49x^8 and simplify
---------------
sqrt(49x^8) = sqrt(49)*sqrt(x^8) = 7*x^4
-----------------------------------------
Cheers,
Stan H.
-----------